Use the following information for the next three questions. Oak Company uses a standard costing system. Manufacturing overhead costs are applied to products on the basis of direct labor hours. The standard cost card shows that 5 direct labor hours at the hourly rate of $25 per hour are required per unit of product. Oak Company had the following data for March: Actual Budgeted Units produced 22,000 20,000 Direct labor hours 105,000 100,000 Variable overhead costs $91,000 $80,000 Fixed overhead costs $51,000 $50,000

1- What was the fixed overhead volume variance for March? A. $ 2,500 Favorable B. $ 2,500 Unfavorable C. $ 5,000 Favorable D. $ 5,000 Unfavorable E. None of the above

2 - Which of the following statements is NOT true?

A. Total manufacturing OH is overapplied by $1,000. B. Variable OH spending variance is $3000 unfavorable. C. Variable OH (Labor) Efficiency variance is $4,000 unfavorable. D. Labor efficiency variance is $125,000 favorable. E. None of the above
Answer #1

Solution 1:

Fixed overhead rate = Budgeted fixed overhead / Budgeted labor hours = $50,000 / 100000 = $0.50 per labor hour

Fixed overhead volume variance = Fixed overhead applied - Budgeted fixed overhead

= (22000*5*$0.50) - $50,000 = $5,000 Favorable

Hence option C is correct.

Solution 2:

Overhead rate = ($80,000 + $50,000) / 100000 = $1.30 per hour

Overhead applied = (22000*5*$1.30) = $143,000

Actual overhead = $91,000 + $51,000 = $142,000

Overapplied overhead = $143,000 - $142,000 = $1,000

Variable overhead rate variance = (SR - AR) * AH = ($0.80 - 91000/105000) * 105000 = $7,000 Unfavorable

Variable overhead efficiency variance = (SH - AH) * SR = (22000*5 - 105000) * $0.80 = $4,000 Favorable

Hence statement that is not true is "Variable OH (Labor) Efficiency variance is $4,000 unfavorable"

Hence option C is correct.